OpenStudy (anonymous):

Commutative Property of Addition OR Associative Property of Addition? 3 + a + b = a + 3 + b

OpenStudy (anonymous):

Can you explain the differences?

Parth (parthkohli):

Commutative property as I suggested. [a + b + c = c + b + a]

Parth (parthkohli):

Associative property is [(a + b) + c = a + (b + c)]

OpenStudy (anonymous):

Thankk you (:

OpenStudy (anonymous):

com.- The order of the plots does not change the sum retrice- The sum does not change when the parcels are associated differently